OpenStudy (aivantettet26):

how many electrons must you have to get a total charge of -1.0c? how many moles of electron is this?

OpenStudy (xishem):

The charge of one electron is:\[-e=-1.602\times 10^{-19}C\]So how would you find how many electrons you need to produce 1C?

OpenStudy (aivantettet26):

\[-6.24x10^{-20}\]

OpenStudy (aivantettet26):

im not really sure :/

OpenStudy (xishem):

Well if one electron has that charge, then N electrons have -1.0C of charge:\[N*(-e)=-1.0C\]\[N(-1.602\times 10^{-19}C)=-1.0C\]And then you can simply solve for N.